INFRARED FLUORESCENCE MEASUREMENTS OF GASEOUS BENZENE WITH A NEW HOME-MADE SPECTROMETER

\underline{G. FÉRAUD}, Y. CARPENTIER^{a}, T. PINO, P. PARNEIX, T. CHAMAILLÉ, Institut des Sciences Moléculaires d’Orsay, Université Paris-Sud 11, Orsay, France; E. DARTOIS, Y. LONGVAL, Institut d’Astrophysique Spatiale, Université Paris-Sud 11, Orsay, France; R. VASQUEZ and Ph. BRÉCHIGNAC, Institut des Sciences Moléculaires d’Orsay, Université Paris-Sud 11, Orsay, France.

Production and characterization of organic molecules such as gas phase Polycyclic Aromatic Hydrocarbons (PAHs) are one of the major challenges in laboratory astrophysics. Infrared spectral signatures are emitted by molecules in interstellar clouds submitted to UV radiations from surrounding stars. These infrared bands, the so-called Aromatic Infrared Bands (AIBs), have been attributed to PAHs since now 27 years^{b}. However, experimental works on infrared emission are quite rare. That is why we are developing an unique infrared spectrometer which collects and detects the infrared emission light from gas phase UV-excited PAHs. We are currently testing the spectrometer with benzene and its derivatives. Their infrared fluorescence decays will be presented, with an attempt to understand the competitive relaxation pathways.

This work was in part supported by the French INSU-CNRS national program ’Physique et Chimie du Milieu Interstellaire (PCMI)’.

aPresent address : Laboratory Astrophysics Group of the Max Planck Institute for Astronomy at the Friedrich Schiller University Jena, Institute of Solid State Physics, Helmholtzweg 3, D-07743 Jena, Germany

bA. Léger and J.-L. Puget, ”Identification of the ’unidentified’ IR emission features of interstellar dust?”, A&A 137(1) :L5-L8